OpenStudy (anonymous):

find all real solutions, leave the answer as a simplifies square root!!! a^2-405=0

geerky42 (geerky42):

Add both sides by 405 then take square root.

geerky42 (geerky42):

Don't forgot \(\pm\)

OpenStudy (anonymous):

I gott hat far and my answer was \[a=+-\sqrt{405}\] an that be sinplified?

OpenStudy (anonymous):

No, you can simplify \[\sqrt{405}\]

OpenStudy (anonymous):

thats where i get stuck.

OpenStudy (anonymous):

9\[\sqrt{5}\]

OpenStudy (anonymous):

\[9\sqrt{5}\]

geerky42 (geerky42):

HINT: 405 = 9·45

geerky42 (geerky42):

= 9·9·5

geerky42 (geerky42):

So √405 = √(9²5) = 9√5

OpenStudy (anonymous):

OH ok \[\sqrt{81}\times \sqrt{5}=9\sqrt{5}???\]

geerky42 (geerky42):

Yeah

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

sweet thanks

geerky42 (geerky42):

You're welcome.